OpenStudy (anonymous):

If u is a vector of length 3, v is a unite vector, and uv=-2, find (u+4v)(v-u)

OpenStudy (anonymous):

uv-uu+4vv-4uv = -uu+ 4vv - 3uv = -9+4-(3*-2) = 1